# AutoAct: Automatic Agent Learning from Scratch via Self-Planning

### v2

@@ Abstract
Language agents have achieved considerable performance on various complex tasks. Despite the incessant exploration in this field, existing language agent systems still struggle with costly, non-reproducible data reliance and face the challenge of compelling a single model for multiple functions. To this end, we introduce AutoAct, an automatic agent learning framework that does not rely on large-scale annotated data and synthetic trajectories from closed-source models (e.g., GPT-4). Given limited data with a tool library, AutoAct first automatically synthesizes planning trajectories without any assistance from humans or strong closed-source models. Then, AutoAct leverages a division-of-labor strategy to automatically differentiate based on the target task information and synthesized trajectories, producing a sub-agent group to complete the task. We conduct comprehensive experiments with different LLMs, which demonstrates that AutoAct yields better or parallel performance compared to various strong baselines. We even notice that AutoAct, when using the Llama-2-13b model, can achieve performance comparable to that of the zero-shot GPT-3.5-Turbo agent [^1] Code will be available at https://github.com/zjunlp/AutoAct..
@@ 1 Introduction
Language agents Wang et al. (2023b); Xi et al. (2023); Xie et al. (2023), which leverage the powerful reasoning capabilities Qiao et al. (2023b); Zhang et al. (2023b) of Large Language Models (LLMs) to generate executable actions for observing the external world, have emerged as essential components of AI systems designed to address intricate interactive tasks Torantulino (2023); Osika (2023); Nakajima (2023); Tang et al. (2023). The process of endowing LLMs with such interactive capabilities is referred to as Agent Learning wherein planning plays a pivotal role, which is responsible for decomposing complex tasks Wei et al. (2022); Zhou et al. (2023a); Yao et al. (2023), invoking external tools Shen et al. (2023); Lu et al. (2023); Qin et al. (2023), reflecting on past mistakes Shinn et al. (2023); Madaan et al. (2023), and aggregating information from various sources to reach the final targets. There have been a lot of works Yao et al. (2023); Wang et al. (2023a); Shen et al. (2023); Chen et al. (2023c) that directly prompt closed-source off-the-shelf LLMs to plan on particular tasks. Despite their convenience and flexibility, closed-source LLMs inevitably suffer from unresolved issues, as their accessibility often comes at a steep price and their black-box nature makes the result reproduction difficult. In light of this, some recent endeavors have shifted their focus towards imbuing open-source models with planning capabilities through fine-tuning Chen et al. (2023a); Zeng et al. (2023); Yin et al. (2023).
@@
  ReAct Yao et al. (2023) User Prompt Iterative
Reflexion Shinn et al. (2023) User Prompt Iterative
Chameleon Lu et al. (2023) User Prompt Global
HuggingGPT Shen et al. (2023) User Prompt Global
AutoGPT Torantulino (2023) User Prompt Iterative
BOLAA Liu et al. (2023) User Prompt Iterative
AgentVerse Chen et al. (2023c) User Prompt Iterative
Agents Zhou et al. (2023c) User Prompt Iterative
AgentTuning Zeng et al. (2023) Benchmark GPT-4 Iterative
FireAct Chen et al. (2023a) Benchmark GPT-4 Iterative
AutoAct (ours) User + Self-Instruct Self-Planning Iterative
However, despite the existing fine-tuning-based method’s achievements, they are not without their limitations. On the one hand, training open-source models necessitates a substantial amount of annotated task data and still relies on closed-source models to synthesize planning trajectories. However, fulfilling these requirements in many real-world scenarios, such as private personal bot or sensitive company business, often proves to be rocky. On the other hand, from the perspective of agent framework design, fine-tuning-based methods compel one single language agent to learn all planning abilities, placing even greater pressure on them. These contradict the Simon’s principle of bounded rationality Mintrom (2015), which states that “precise social division-of-labor and clear individual tasks can compensate for the limited ability of individuals to process and utilize information”.
To this end, we introduce AutoAct, an automatic agent learning framework, which does not rely on large-scale annotated data and synthetic trajectories from closed-source models while incorporating explicit individual tasks with precise division-of-labor (see Figure 1). Given a limited set of user-provided data examples, AutoAct starts with a Meta-Agent to obtain an augmented database through self-instruct Wang et al. (2023c). Then, armed with a prepared tool library, the Meta-Agent can automatically synthesize planning trajectories without any assistance from humans or strong closed-source models. Finally, we propose the division-of-labor strategy which resembles cell differentiation based on the self-synthesized trajectories (genes), where the Meta-Agent acts as a stem cell Colman (2008) and differentiates into three sub-agents with distinct functions: task decomposition, tool invocation, and self-reflection, respectively. Our differentiation process is essentially a parameter-efficient training process on the self-synthesized trajectories with low consumption resources. We list the differences between AutoAct and prior works in Table 1.
Experiments on complex question-answering tasks with different LLMs demonstrate that the proposed AutoAct yields better or parallel performance compared to various strong baselines. We summarize our main contributions as follows:
  • We propose AutoAct, an automatic agent learning framework that does not rely on large-scale annotated data and synthetic trajectories from closed-source models while adhering to the principle of bounded rationality.
  • We conduct comprehensive experiments with different LLMs, which demonstrates that AutoAct yields better or parallel performance compared to various strong baselines. We even notice that when using the Llama-2-13b model, AutoAct can achieve performance comparable to that of the zero-shot GPT-3.5-Turbo agent OpenAI (2022).
  • Extensive empirical analysis demonstrates the effectiveness of our appropriate division-of-labor strategy and the trajectory quality generated by AutoAct outperforms that of other methods from multiple aspects.
@@ 2 AutoAct
Overview.
As shown in Figure 2, AutoAct only requires target task information and a language agent (we name it Meta-Agent) to initiate its work. The Meta-Agent first augments the task data from scratch by self-instruct. Furthermore, with a tool library available, the Meta-Agent conducts automatic agent learning by differentiating into sub-agents with distinct functionalities and enabling them to perform group task-specific planning. We name this process as self-planning. Below is a detailed introduction of AutoAct. Note that all symbols used are globally defined.
@@ 2.1 Critical Components of AutoAct
The Meta-Agent stands at the central position of our AutoAct framework. It is responsible for all the preparatory work before self-differentiation and serves as the backbone model for the self-differentiation process. Given limited target task information and a pre-prepared tool library, the Meta-Agent can differentiate into an agent group capable of collaborating to accomplish the target task. In AutoAct, the Meta-Agent can be initialized with any open-source model.
@@
In this paper, we mainly focus on agent learning from scratch, which means the task information at hand is quite limited, primarily encompassing three aspects: task name $\mathcal{M}$, task description $\mathcal{P}$, task data examples $\mathcal{C}$. Concretely, $\mathcal{P}$ represents a detailed description of the task’s characteristics, properties, and other relevant information. $\mathcal{C}=\{q_{i},a_{i}\}_{i=1}^{|\mathcal{C}|}$ indicates $|\mathcal{C}|$ question-answer example pairs of the task, where $|\mathcal{C}|$ is very small which users can effortlessly provide (e.g., a few demonstrations). For a more in-depth view of task information, please refer to Appendix B. Note that the task information serves as the only user-provided knowledge of the task for AutoAct to conduct automatic agent learning.
@@
To facilitate our agents in automatic task planning, we provide a comprehensive tool library at their disposal. The tool library can be denoted as $\mathcal{T}=\{m_{i},d_{i},u_{i}\}_{i=1}^{|\mathcal{T}|}$, where $m$ represents the name of each tool, $d$ defines the functionality of each tool, $u$ details the usage instruction of each tool, and $|\mathcal{T}|$ stands for the tools amount of the library. In our automatic procedure, the Meta-Agent has the autonomy to select appropriate tools from the tool library based on the task information. Users also have the option to expand the tool library according to their specific needs, allowing for more flexible utilization. We list part of our tool library in Appendix C.
@@ 2.2 Starting from Scratch via Self-Instruct
To acquire a sufficient amount of task data and provide an ample training resource, it is necessary to augment the data based on the examples at hand. We accomplish this process through self-instruct. Initially, the database $\mathcal{D}$ is set to be equal to the task data examples $\mathcal{C}$, with $\mathcal{C}$ as the seed for data generation. In each round, the Meta-Agent generates new question-answer pairs by few-shot prompting, and the few-shot prompt examples are randomly sampled from $\mathcal{D}$. The generated data will be added to $\mathcal{D}$ followed by filtering, with the exclusion of format erroneous and duplicate data before its inclusion. Eventually, we obtain a database $\mathcal{D}=\{q_{i},a_{i}\}_{i=1}^{|\mathcal{D}|}$, where the number of data $|\mathcal{D}|$ satisfies $|\mathcal{D}|\gg|\mathcal{C}|$. The prompt we use for self-instruct can be seen in Appendix D.1.
@@ 2.3 Automatic Agent Learning via Self-Planning
With the tool library at hand, we ask the Meta-Agent to select applicable tools for each task automatically. Specifically, we put $\mathcal{T}=\{m_{i},d_{i},u_{i}\}_{i=1}^{|\mathcal{T}|}$ in the form of a tool list as part of the prompt. Along with $\mathcal{T}$, the prompt also includes the task’s description $\mathcal{C}$. Finally, we instruct the Meta-Agent to select an appropriate set of tools $\mathcal{T}_{s}$ ($\mathcal{T}_{s}\subset\mathcal{T}$) to wait for synthesizing trajectories. The prompt we use for automatic tool selection can be seen in Appendix D.2.
@@
Without relying on closed-source models, we enable the Meta-Agent to synthesize planning trajectories on its own. Equipped with $\mathcal{T}_{s}$, we instruct the Meta-Agent to synthesize trajectories in a zero-shot manner on the database $\mathcal{D}$ adhering to the format of Thought-Action-Observation as defined in Yao et al. (2023). In order to obtain high-quality synthesized trajectories, we filter out all the trajectories with $\texttt{reward}<1$ and collect trajectories with exactly correct answers ($\texttt{reward}=1$) as the training source for self-differentiation. The prompt for trajectories synthesis can be seen in Appendix D.3.
@@
    Plan-Agent $\pi_{\rm plan}$ undertakes task decomposition and determines which tool to invoke in each planning loop (Eq.2).
@@
where $\tau^{r}$ and $\alpha^{r}$ represent the thought and action of the reflection process respectively, and $\mathcal{H}$ is the planning history after finishing the answer. The trajectories can be reorganized based on the responsibilities above and fed to the Meta-Agent for self-differentiation. Our differentiation is a parameter-efficient fine-tuning process to achieve resource-efficient learning. Particularly, for each sub-agent, we train a specific LoRA Hu et al. (2022).
@@
After obtaining the task-specific sub-agents, any new question is processed through group planning among the sub-agents to achieve the desired outcome. Once the tool name $\alpha_{t}^{m}$ generated by the Plan-Agent is triggered at time $t$, the Tool-Agent is roused to determine the parameters $\alpha_{t}^{p}$ transferred to the specific tool. The return result of the tool is treated as the observation $o_{t}$ and handed to the Plan-Agent. After the collaboration between the Plan-Agent and Tool-Agent finishes with a prediction, the Reflect-Agent comes on the stage to reflect on the history and provide a reflection result contained in the reflection action $\alpha^{r}$. If the reflection result indicates that the prediction is correct, the whole planning process concludes. Otherwise, the Plan-Agent and Tool-Agent will continue the planning based on the reflection information. The specific sequence of the group planning process can be referred to the trajectory example on the right side of Figure 2.
@@
3 Experimental Setup
Tasks.
We evaluate our method on HotpotQA Yang et al. (2018) and ScienceQA Lu et al. (2022). HotpotQA is a multi-hop QA task challenging for rich background knowledge, the answer of which is usually a short entity or yes/no. Following Liu et al. (2023), we randomly select 300 dev questions divided into three levels for evaluation, with 100 questions in each level. For HotpotQA, the $\texttt{reward}\in[0,1]$ is defined as the F1 score grading between the prediction and ground-truth answer. ScienceQA is a multi-modal QA task spanning various scientific topics. We also divide the test set into three levels based on the grade, with 120 randomly sampled data in each level. Since ScienceQA is a multi-choice QA task, the $\texttt{reward}\in\{0,1\}$ is exactly the accuracy. Note that due to the limitations of LMs in generating images, for ScienceQA, during the self-instruct stage, we directly generate captions for the images instead.
Baselines.
We choose the open-source Llama-2 models Touvron et al. (2023) as the backbones of our Meta-Agent and sub-agents. The compared baselines are as follows: 1) CoT Wei et al. (2022), the naive Chain-of-Thought reasoning method. 2) ReAct Yao et al. (2023), a well-known single-agent framework based on few-shot learning that performs planning and action iteratively. 3) Chameleon Lu et al. (2023), another few-shot single-agent framework that performs planning before action. 4) Reflexion Shinn et al. (2023), a single-agent framework to reinforce language agents through linguistic feedback. 5) BOLAA Liu et al. (2023), a multi-agent framework that customizes different agents through prompts. 6) FireAct Chen et al. (2023a), a single-agent framework with fine-tuning on diverse kinds of trajectories generated by GPT-4 OpenAI (2023). 7) GPT-3.5-Turbo OpenAI (2022). To ensure fairness, we maintain an equal training trajectory volume of 200 for FireAct and AutoAct (200 synthesized data). As Reflexion provides answer correctness labels during reflection but other methods including AutoAct do not, we test all the other methods twice and choose the correct one for evaluation. For all the prompt-based baselines, we uniformly provide two examples in the prompt.
Training Setups.
We fine-tune all our models with LoRA Hu et al. (2022) in the format proposed in Alpaca Taori et al. (2023). Our fine-tuning framework leverages FastChat Zheng et al. (2023) using DeepSpeed Rasley et al. (2020). We set the learning rate of 1e-4 and the sequence length of 4096 for all the model scales. The training epoch for the 13b and 7b models is 5 and for the 70b model is 3. The batch size is 4 for the 13b and 7b models and 1 for the 70b model. We use the AdamW optimizer Loshchilov and Hutter (2019) with a cosine learning scheduler. We detail the hyper-parameters for training in Appendix A.
4 Results
Compare to Prompt-based Agent Learning Baselines.
As shown in Table 2, the 13b and 70b models consistently outperform various prompt-based baselines. The 70b model even surpasses the agent performance of GPT-3.5-Turbo, achieving a rise of $\uparrow$3.77on HotpotQA and $\uparrow$6.39on ScienceQA. The performance of the 7b model is comparable to other methods to some extent. However, it fails to deliver the same impressive results due to its limited self-instruct and self-planning capabilities. Therefore, whether in a single-agent or multi-agent architecture, prompt-based methods relying on few-shot demonstrations fail to precisely customize the behavior of the agent, which is also supported by the fact that FireAct widely outperforms ReAct and BOLAA in the context of iterative planning. In addition, our investigation reveals a visible disparity in open-source models between the performance of many prompt-based planning baselines (relying on various external tools) and CoT (relying on the models’ intrinsic reasoning abilities). This discrepancy underscores the formidable challenge of unlocking planning capabilities by prompting.
Compare to Fine-tuning-based Agent Learning Baselines.
Further focusing on FireAct in Table 2, despite the assistance of GPT-4, FireAct’s approach of assigning the entire planning task to a single model proves to be burdensome. As a result, its performance on ScienceQA even falls short compared to the prompt-based global planning method, Chameleon. AutoAct employs self-differentiation to decouple the planning process and reaches a clear division-of-labor among sub-agents for group planning, resulting in a improvement than FireAct, with an enhancement of $\uparrow$5.77on HotpotQA and $\uparrow$6.67on ScienceQA with 70b model. Additionally, AutoAct achieves self-planning without relying on closed-source models and large-scale labeled datasets, which paves the way for automatic agent learning with open-source models from scratch. In ablation study (§4) and human evaluation (§5), we will further validate that the quality of trajectories synthesized by AutoAct is not inferior to FireAct trained on trajectories synthesized using GPT-4.
Figure 3: Performance of AutoAct on different training data scales. (a-c) represents the results of the model trained on self-synthesized trajectories. (d-f) represents the results of the model trained on trajectories synthesized by a stronger model, where the dashed line is the baseline trained on self-synthesized trajectories.
Figure 4: Performance of AutoAct based on different degrees of labor division. One is training a single model with all the differentiated data. Three represents the differentiation into three agents: plan, tool, and reflect. Tool Specified indicates further differentiating the tool-agent with one tool, one agent.
HotpotQA ScienceQA
  AutoAct 48.47 78.61

### v4

# AutoAct: Automatic Agent Learning from Scratch for QA via Self-Planning

@@ Abstract
Language agents have achieved considerable performance on various complex question-answering tasks by planning with external tools. Despite the incessant exploration in this field, existing language agent systems still struggle with costly, non-reproducible data reliance and face the challenge of compelling a single model for multiple functions. To this end, we introduce AutoAct, an automatic agent learning framework for QA that does not rely on large-scale annotated data and synthetic planning trajectories from closed-source models (e.g., GPT-4). Given limited data with a tool library, AutoAct first automatically synthesizes planning trajectories without any assistance from humans or strong closed-source models. Then, AutoAct leverages a division-of-labor strategy to automatically differentiate based on the target task information and synthesized trajectories, producing a sub-agent group to complete the task. We conduct comprehensive experiments with different LLMs, which demonstrates that AutoAct yields better or parallel performance compared to various strong baselines. Further analysis demonstrates the effectiveness of the division-of-labor strategy, with the trajectory quality generated by AutoAct generally outperforming that of others [^1] Code: https://github.com/zjunlp/AutoAct..
@@ 1 Introduction
Language agents Wang et al. (2023a); Xi et al. (2023); Guo et al. (2024), which leverage the powerful reasoning capabilities Qiao et al. (2023b); Zhang et al. (2023) of Large Language Models (LLMs) to interact with executable tools, have emerged as essential components of AI systems designed to address complex question-answering tasks Torantulino (2023); Osika (2023); Nakajima (2023); Tang et al. (2023); Xie et al. (2023). The process of endowing LLMs with such interactive capabilities is referred to as Agent Learning wherein planning Huang et al. (2024b) plays a pivotal role, which is responsible for decomposing complex questions into simpler ones Wei et al. (2022); Yao et al. (2023); Team (2023); Qian et al. (2023), invoking external tools Shen et al. (2023); Lu et al. (2023); Qin et al. (2023), reflecting on past mistakes Shinn et al. (2023); Madaan et al. (2023), and aggregating information from various sources to reach the final answer. There have been a lot of works Li et al. (2023); Shen et al. (2023); Hong et al. (2023); Talebirad and Nadiri (2023); Chen et al. (2023d, b) that directly prompt closed-source off-the-shelf LLMs to plan on particular tasks. Despite their convenience and flexibility, closed-source LLMs inevitably suffer from unresolved issues, as their accessibility often comes at a steep price and their black-box nature makes the result reproduction difficult. In light of this, some recent endeavors have shifted their focus towards imbuing open-source models with planning capabilities through fine-tuning Chen et al. (2023a); Zeng et al. (2023); Yin et al. (2023).
@@
However, despite the achievements of the existing fine-tuning-based methods, they are not without limitations. On the one hand, training open-source models necessitates a substantial amount of annotated QA data pairs and still relies on closed-source models to synthesize planning trajectories. However, fulfilling these requirements in many real-world scenarios, such as private personal bots or sensitive company business, often proves to be rocky. On the other hand, from the perspective of agent framework, fine-tuning-based methods compel one single language agent to learn all planning abilities, placing even greater pressure on them. These contradict Simon’s principle of bounded rationality (Mintrom, 2015), which states that “precise social division-of-labor and clear individual tasks can compensate for the limited ability of individuals to process and utilize information”.
To this end, we introduce AutoAct, an automatic agent learning framework for QA, which does not rely on large-scale annotated data and synthetic trajectories from closed-source models while incorporating explicit individual tasks with precise division-of-labor (see Fig. 1). Given a limited set of user-provided data examples, AutoAct starts with a Meta-Agent to obtain an augmented database through self-instruct (Wang et al., 2023b). Then, armed with a prepared tool library, the Meta-Agent can automatically synthesize planning trajectories without any assistance from humans or strong closed-source models. Finally, we propose the division-of-labor strategy which resembles cell differentiation based on the self-synthesized trajectories (genes), where the Meta-Agent acts as a stem cell (Colman, 2008) and differentiates into three sub-agents with distinct functions: task decomposition, tool invocation, and self-reflection, respectively. Our differentiation process is essentially a parameter-efficient training process on the self-synthesized trajectories with low-consumption resources. We list the differences between AutoAct and prior works in Tab. 3.
Experiments on complex question-answering tasks with different LLMs demonstrate that AutoAct yields better or parallel performance compared to various strong baselines. Extensive empirical analysis demonstrates the effectiveness of our appropriate division-of-labor strategy.
@@ 2.1 Critical Components of AutoAct
The Meta-Agent is responsible for all the preparatory work before self-differentiation and serves as the backbone model for all sub-agents. Given limited target task information and a pre-prepared tool library, the Meta-Agent can differentiate into an agent group capable of collaborating to accomplish the target task. In AutoAct, the Meta-Agent can be initialized with any kind of open-source model.
@@
In this paper, we mainly focus on agent learning from scratch, which means the task information at hand is quite limited, primarily encompassing three aspects: task name $\mathcal{M}$, task description $\mathcal{P}$, task data examples $\mathcal{C}$. Concretely, $\mathcal{P}$ represents a detailed description of the task’s characteristics. $\mathcal{C}=\{q_{i},a_{i}\}_{i=1}^{|\mathcal{C}|}$ indicates $|\mathcal{C}|$ question-answer example pairs of the task, where $|\mathcal{C}|$ is very small which users can effortlessly provide (e.g., a few demonstrations). For a more in-depth view of task information, please refer to Appx. E. Note that the task information serves as the only user-provided knowledge of the task for AutoAct to conduct automatic agent learning.
@@
To facilitate our agents in automatic task planning, we provide a comprehensive tool library at their disposal. The tool library can be denoted as $\mathcal{T}=\{m_{i},d_{i},u_{i}\}_{i=1}^{|\mathcal{T}|}$, where $m$ represents the tool name, $d$ defines the tool functionality, $u$ details the tool usage instruction, and $|\mathcal{T}|$ stands for the tool amount of the library. In our automatic procedure, the Meta-Agent has the autonomy to select appropriate tools from the tool library based on the task information. Users also have the option to expand the tool library according to their specific needs, allowing for more flexible utilization. We list the details of our tool library in Appx. F.
@@ 2.2 Starting from Scratch via Self-Instruct
To acquire a sufficient amount of task data and provide an ample training resource, it is necessary to augment the data based on the examples at hand. We accomplish this process through self-instruct. Initially, the database $\mathcal{D}$ is set to be equal to the task data examples $\mathcal{C}$, with $\mathcal{C}$ as the seed for data generation. In each round, the Meta-Agent generates new question-answer pairs by few-shot prompting, and the few-shot prompt examples are randomly sampled from $\mathcal{D}$. The generated data will be added to $\mathcal{D}$ followed by filtering, with the exclusion of format erroneous and duplicate data before its inclusion. Eventually, we obtain a database $\mathcal{D}=\{q_{i},a_{i}\}_{i=1}^{|\mathcal{D}|}$, where the number of data $|\mathcal{D}|$ satisfies $|\mathcal{D}|\gg|\mathcal{C}|$. The prompt we use for self-instruct can be seen in Appx. G.1 and we list some cases generated through self-instruct in Appx. H.
@@ 2.3 Automatic Agent Learning via Self-Planning
With the tool library at hand, we ask the Meta-Agent to select applicable tools for each task automatically. Specifically, we put $\mathcal{T}=\{m_{i},d_{i},u_{i}\}_{i=1}^{|\mathcal{T}|}$ in the form of a tool list as part of the prompt. Along with $\mathcal{T}$, the prompt also includes the task’s description $\mathcal{C}$. Finally, we instruct the Meta-Agent to select an appropriate set of tools $\mathcal{T}_{s}$ ($\mathcal{T}_{s}\subset\mathcal{T}$) to wait for synthesizing trajectories. The prompt we use for automatic tool selection can be seen in Appx. G.2.
@@
Without depending on closed-source models, we enable the Meta-Agent to synthesize planning trajectories on its own. Equipped with $\mathcal{T}_{s}$, we instruct the Meta-Agent to synthesize trajectories in a zero-shot manner on the database $\mathcal{D}$ adhering to the format of Thought-Action-Observation as defined in Yao et al. (2023). In order to obtain high-quality synthesized trajectories, we filter out all the trajectories with $\texttt{reward}<1$ and collect trajectories with exactly correct answers ($\texttt{reward}=1$) as the training source for self-differentiation. The prompt for trajectories synthesis can be seen in Appx. G.3.
@@
    Plan-Agent $\pi_{\rm plan}$ undertakes question decomposition and determines which tool to invoke in each planning loop (Eq. 2).
@@
where $\tau^{r}$ and $\alpha^{r}$ represent the thought and action of the reflection process, and $\mathcal{H}$ is the planning history after finishing the answer. The trajectories can be reorganized based on the responsibilities above and fed to the Meta-Agent for self-differentiation. Our differentiation is a parameter-efficient fine-tuning process to achieve resource-efficient learning. We give examples of the training data for each sub-agent in Appx. I. Particularly, for each sub-agent, we train a specific LoRA Hu et al. (2022).
@@
At inference time, once the tool name $\alpha_{t}^{m}$ generated by the Plan-Agent is triggered at time $t$, the Tool-Agent is roused to decide the parameters $\alpha_{t}^{p}$ transferred to the specific tool. The return result of the tool is treated as the observation $o_{t}$ and handed to the Plan-Agent. After the collaboration between the Plan-Agent and Tool-Agent reaches a prediction, the Reflect-Agent comes to reflect on the history and provide a reflection result contained in the reflection action $\alpha^{r}$. If the reflection result indicates that the prediction is correct, the whole planning process ends. Otherwise, the Plan-Agent and Tool-Agent will continue the planning based on the reflection information. The specific sequence of the group planning process can be found in the example on the right of Fig. 2.
@@
3 Experimental Setup
Tasks and Metrics.
We evaluate AutoAct on HotpotQA Yang et al. (2018) and ScienceQA Lu et al. (2022). HotpotQA is a multi-hop QA task challenging for rich background knowledge, the answer of which is usually a short entity or yes/no. Following Liu et al. (2023), we randomly select 300 dev questions divided into three levels for evaluation, with 100 questions in each level. For HotpotQA, the $\texttt{reward}\in[0,1]$ is defined as the F1 score grading between the prediction and ground-truth answer. ScienceQA is a multi-modal QA task spanning various scientific topics. We also divide the test set into three levels based on the grade, with 120 randomly sampled data in each level. Since ScienceQA is a multi-choice task, the $\texttt{reward}\in\{0,1\}$ is exactly the accuracy. Note that due to the limitations of LMs in generating images, for ScienceQA, during the self-instruct stage, we directly generate captions for the images instead.
Baselines.
We choose the open-source Llama-2 models Touvron et al. (2023) and Mistral-7B Jiang et al. (2023) as the backbones of our Meta-Agent and sub-agents. The compared baselines include CoT Wei et al. (2022), ReAct, Chameleon Lu et al. (2023), Reflexion Shinn et al. (2023), BOLAA Liu et al. (2023), ReWOO Xu et al. (2023), FireAct Chen et al. (2023a). We detail each baseline in Appx. B. To ensure fairness, we maintain an equal training trajectory volume of 200 for FireAct and AutoAct (200 synthesized data). As Reflexion provides answer correctness labels during reflection but other methods including AutoAct do not, we test all the other methods twice and choose the correct one for evaluation. For all the prompt-based baselines, we uniformly provide two examples in the prompt.
Training Setups.
We fine-tune all our models with LoRA Hu et al. (2022) in the format proposed in Alpaca Taori et al. (2023). All the training and inference experiments are conducted on 8 V100 GPUs within 16 hours. We detail the hyper-parameters for training in Appx. B.
4 Results
Compare to Prompt-based Agent Learning Baselines.